OpenStudy (lgbasallote):

how to get a solution for \(\sqrt{1-y^2} dx + \sqrt{1-x^2}dy = 0\)?

OpenStudy (anonymous):

seperable variables

myininaya (myininaya):

Agreed! :)

OpenStudy (lgbasallote):

is it possible to use linear DE?

OpenStudy (anonymous):

subtract one to one side then divide by the GCF \[-\sqrt{1-y^2}\sqrt{1-x^2}\]

OpenStudy (anonymous):

why would you want to?

OpenStudy (lgbasallote):

because it is a problem in the linear DE section of the book lol

OpenStudy (anonymous):

i thnk that would make it harder than it should. Your book is just tryng to make you be able to distinguish between certain ways of solving DE's

OpenStudy (anonymous):

literally there is numerous of ways you could solve this however seperation by variables would be the most logical way to do this

OpenStudy (lgbasallote):

hmm ok..so how to do with variable separable?

OpenStudy (lgbasallote):

hmm okay..

OpenStudy (anonymous):

my bad it should be \sqrt{1-x^2}dy=-\sqrt{1-y^2}dx\]

OpenStudy (anonymous):

\[\sqrt{1-x^2}dy=-\sqrt{1-y^2}dx\]

OpenStudy (lgbasallote):

then divide?

OpenStudy (anonymous):

then find a greest commonfactor to cancel out each term that is next to the differentials

OpenStudy (lgbasallote):

GCF??

OpenStudy (anonymous):

it'd be basically multiply both sides by 1/both functions

OpenStudy (lgbasallote):

oh so it is divide lol

OpenStudy (anonymous):

i'll show you... it's a way of doing the dividing in one step

OpenStudy (lgbasallote):

okay :S

OpenStudy (anonymous):

\[\frac{1}{\sqrt{1-x^2}\sqrt{1-y^2}}(\sqrt{1-x^2})dy=\frac{1}{\sqrt{1-x^2}\sqrt{1-y^2}}(-\sqrt{1-y^2})dx\]

OpenStudy (anonymous):

see how the comon factor cancels it out on both sides

OpenStudy (anonymous):

you're left with \[\frac{dy}{\sqrt{1-y^2}}=-\frac{dx}{\sqrt{1-x^2}}\]

OpenStudy (lgbasallote):

hmm yep

OpenStudy (anonymous):

now you can integrate both sides

OpenStudy (lgbasallote):

so it's trig sub?

OpenStudy (anonymous):

yep i believe that is arcsin

OpenStudy (lgbasallote):

thanks :D